# Multi-objective Combinatorial Methodology for Nuclear Reactor Site Assessment: A Case Study for the United States

### v1

@@ Abstract
As the global demand for clean energy intensifies to achieve sustainability and net-zero carbon emission goals, nuclear energy stands out as a reliable solution. However, fully harnessing its potential requires overcoming key challenges, such as the high capital costs associated with nuclear power plants (NPPs). One promising strategy to mitigate these costs involves repurposing sites with existing infrastructure, including coal power plant (CPP) locations, which offer pre-built facilities and utilities. Additionally, brownfield sites—previously developed or underutilized lands often impacted by industrial activity—present another compelling alternative. These sites typically feature valuable infrastructure that can significantly reduce the costs of NPP development. This study introduces a novel multi-objective optimization methodology, leveraging combinatorial search to evaluate over 30,000 potential NPP sites in the United States. Our approach addresses gaps in the current practice of assigning pre-determined weights to each site attribute that could lead to bias in the ranking. Each site is assigned a performance-based score, derived from a detailed combinatorial analysis of its site attributes. The methodology generates a comprehensive database comprising site locations (inputs), attributes (outputs), site score (outputs), and the contribution of each attribute to the site score (outputs). We then use this database to train a machine learning neural network model, enabling rapid predictions of nuclear siting suitability across any location in the contiguous United States. By interpolating between the sites in the database, the model addresses critical siting metrics with exceptional efficiency. Our findings highlight that both brownfield and CPP sites are highly competitive for nuclear development. Notably, three brownfield sites in California and three CPP sites in Wisconsin, North Carolina, and Pennsylvania rank among the most promising locations. These results underscore the potential of integrating machine learning and optimization techniques to transform nuclear siting, paving the way for a cost-effective and sustainable energy future.
@@
Nuclear Power Plant, Brownfield, Site Selection, Multi-Objective Optimization, Coal to Nuclear Transition
journal: Elsevier
@@
Feed-Forward Neural Network
@@ 1 Introduction
Nuclear energy has several important applications, ranging from electricity generation to specialized industrial and medical uses [1, 2, 3, 4]. Nuclear energy ranks among the highest-yielding energy sources and has very low carbon dioxide emissions, even when accounting for the entire lifespan of the nuclear power plant (NPP). With its significant potential, nuclear energy could play a key role in achieving net-zero carbon emissions. However, a major challenge is its high levelized cost of electricity. While production costs for renewable energy sources have consistently declined over time, nuclear energy had the highest levelized cost of electricity in 2023, driven by rising capital expenditures and regulatory requirements [5]. In contrast, wind and solar energy tend to reduce electricity prices in their regions since their levelized costs are generally lower than the average for all energy sources [6]. Although the raw material costs for nuclear electricity production have also decreased [7], the substantial initial costs of nuclear power facilities remain a primary factor in nuclear energy’s higher prices and contribute to public concerns and negative sentiment toward nuclear power along with other issues, including radioactive waste and safety [8, 9]. The construction of supporting infrastructure is one of the expensive prerequisites for nuclear power.
@@
By repurposing established facilities, the capital costs associated with NPP construction could be reduced. This analysis focuses on the United States, where we consider two existing facility options for new nuclear siting: coal power plant (CPP)s and brownfields. First, we consider using the CPP station sites for NPP development. In comparison to a greenfield NPP building operation, the overall capital costs could be decreased by 15-35%, depending on the magnitude and choice of technology used in repurposing the framework of a coal power facility with a 1,200 MWe power output [10]. However, these CPP locations still have both drawbacks. For example, if site remediation of a CPP is required before beginning the NPP construction project, the coal to nuclear repurposing costs could be high. Still, the existing educated workforce of the CPP would decrease the training and education costs of the personnel [11]. A complete cost analysis must be completed before such a repurposing operation. According to the U.S. Department of Energy, 237 operational and 157 idle potential CPP sites are appropriate for hosting nuclear energy facilities [10]. Before reusing the CPP buildings, each of these sites must be thoroughly studied, as NPPs are complex, long-term projects with several risk factors and consequences. The NPP sitting in U.S. CPP locations has been researched in the past [12, 13, 14, 15]. CPP locations promise to be valuable sites for NPP siting. However, the number of U.S. CPPs is not enough to achieve the ultimate net-zero goals for the country. Therefore, an analysis that includes a more comprehensive look at all potential NPP sites in the U.S. is needed.
@@
The second option to host a NPP is to consider brownfield sites. The term ”brownfield” refers to formerly built land that may be contaminated, underutilized, or abandoned, and may occasionally have low pollution levels [16]. The economic costs and risks of brownfield redevelopment projects are lower in areas that have an increased demand for residential and commercial development [17]. According to the U.S. Environmental Protection Agency (EPA), Assessment Cleanup and Redevelopment Exchange System (ACRES), brownfield sites are properties where the existence or potential presence of a hazardous material, pollutant, or contaminant may make it more difficult to expand, redevelop, or reuse [18]. Cmmon types of brownfield sites are former industrial or manufacturing plants, abandoned gas stations, landfills or waste disposal areas, rail yards, mines, and junkyards. Repurposing these sites could be beneficial for the environment and economy. To assess, clean up, and sustainably reuse the brownfield sites, The EPA Brownfields Program provides grants and technical assistance to communities.
@@
The siting problem of NPP location has been an ongoing topic for the last 80 years. The problem consists of many different degrees of freedom and each requirement must be analyzed carefully since NPPs are long-term and high-cost projects. Most of the siting analyses in this field use a technique called the “weighted sum method”. In this method, subjective weights are applied to the geographical data to perform location selection studies for NPPs. To form an objective function, these techniques use pre-determined weights that the analyst justifies based on several factors. Geographical characteristics like seismicity, the presence of cooling water, fault lines, the distance to centers of population, international borders, and logistic network are the most crucial factors taken into account in this weighted sum method [13]. In some studies, only the most crucial characteristics are taken into account, such as seismic activity, population density, cooling water cost, reactor unit cost, and consumer proximity [14]. More sophisticated approaches use detailed NPP socioeconomic data. These methods handle financial, site-related, welfare, and project life-cycle factors independently in separate subroutines, by using their built-in socioeconomic data and weights [15]. Both simple and complex approaches still depend on users to assess the weights assigned to each site objective or attribute. The analyst responsible for setting thresholds, assigning weights, and processing raw data ultimately makes these critical decisions. In this process, the choice of weights can be subjective or objective, but never entirely removes the analyst’s bias in prioritizing certain sites for location selection. This is demonstrated in a recent study on CPP assessment in the U.S. for NPP transition conducted by [19], which involved assigning weights to various site objectives. While we recognize that the authors determined these weights through thorough research and extensive consultations, this level of rigor is cannot be guaranteed in every study. Additionally, the approach still introduces bias, as site rankings inherently depend on the assigned weights. Therefore, an alternative method that avoids predetermined weights is needed to better identify ideal locations for NPP siting with limited bias. Multi-objective optimization techniques may offer some solution pathways to these challenges.
@@
A range of optimization methods for minimization and maximization problems are available in the literature, typically grouped into four main categories: stochastic, genetic, first-order, and second-order algorithms [20]. Most optimization algorithms can solve single-objective optimization problems when the objective orfitness or cost function is a scalar. However, when addressing problems that involve multiple objectives, multi-objective optimization (MOO) approaches are required. These algorithms are designed to handle the complexity of balancing multiple objectives simultaneously during the search.
@@
In the problem of site selection, there is a high likelihood that multiple sites may have favorable characteristics which makes optimization challenging. We propose the concept of site characteristic “domination” as a solution to this problem. The method of non-dominated sorting is used in different optimization algorithms to identify optimal solutions in [21, 22, 23]. The most optimal samples in a non-dominated sorting operation are also known as the Pareto front, where each solution represents a trade-off across multiple objectives. In this Pareto-efficient set, no single solution is superior across all objectives, but the Pareto solutions that balance all objectives can be further analyzed. Consequently, in order to find the best NPP siting locations independent of any bias, we conduct a combinatorial analysis of Pareto front solutions by running a massive number of combinations of site attributes, thereby mitigating external bias towards any site.
This research presents a new approach to nuclear reactor site assessment and brownfield redevelopment by specifically targeting the siting of nuclear power plants NPPs; an area that has not been thoroughly explored in the existing literature. Utilizing raw site data available through the STAND tool, we created a dataset of brownfield NPP siting locations, focusing on 34,548 sites that do not require extensive industrial cleanup, thereby establishing a robust resource for future nuclear siting projects. Furthermore, we conducted a comparative analysis between CPPs and brownfields, challenging the prevailing assumption that CPP sites are the only sites that offer optimal conditions for NPP siting. By comparing 265 CPPs with the brownfield sites, our findings illuminate the feasibility of alternative sites, thereby offering additional compelling options to the current trends in coal-to-nuclear transition research. To enhance the efficiency of future site assessment, we developed a machine learning model that predicts siting metrics and objective importance; significantly reducing the computational burden associated with traditional methodologies. This model is poised to facilitate rapid assessments for any contiguous U.S. location, ultimately streamlining the decision-making process in NPP siting. Accordingly, the key contributions of this study can be summarized as:
@@ 2 Data Collection and Processing
In the brownfield nuclear reactor siting analysis, we use the ACRES brownfield database (described above) as a starting point. We gathered the geographical database files from the EPA. We extracted the coordinates recorded in ACRES from the database layers. We then use the extracted coordinates in the Siting Tool for Advanced Nuclear Development (STAND) tool developed by Fastest Path to Zero Initiative (FPtZ) at the University of Michigan [24]. We processed each extracted brownfield coordinate STAND to creating a brownfield nuclear reactor siting dataset that includes all brownfield sites with their corresponding coordinates and site characteristics. We found unreliable data for Alaska and Hawaii in the STAND tool,because of data availability issues; the brownfield sites in these two states were excluded from the final dataset. The ACRES U.S. brownfield locations gathered from EPA are given in Figure 1. The brownfield siting dataset is current as of May 2024.
About 34,212 different contiguous U.S. brownfield coordinates in the EPA ACRES brownfield database have been processed through STAND to generate values for 22 unique objectives. The attributes in the dataset describe the characteristics of each site. We chose these objectives the 1000 MWt or higher power reactor siting parameters used in the coal-to-nuclear transition analysis performed by the previous NPP siting research [19]. We classify the objectives selected for the brownfield sites under three categories:
@@
Figure 1: ACRES brownfield locations provided by EPA.

### v2

@@ Abstract
As clean energy demand grows to meet sustainability and net-zero goals, nuclear energy emerges as a reliable option. However, high capital costs remain a challenge for nuclear power plants (NPP), where repurposing coal power plant sites (CPP) with existing infrastructure is one way to reduce these costs. Additionally, Brownfield sites—previously developed or underutilized lands often impacted by industrial activity—present another compelling alternative. This study introduces a novel multi-objective optimization methodology, leveraging combinatorial search to evaluate over 30,000 potential NPP sites in the United States. Our approach addresses gaps in the current practice of assigning pre-determined weights to each site attribute that could lead to bias in the ranking. Each site is assigned a performance-based score, derived from a detailed combinatorial analysis of its site attributes. The methodology generates a comprehensive database comprising site locations (inputs), attributes (outputs), site score (outputs), and the contribution of each attribute to the site score. We then use this database to train a neural network model, enabling rapid predictions of nuclear siting suitability across any location in the United States. Our findings highlight that CPP sites are highly competitive for nuclear development, but some Brownfield sites are able to compete with them. Notably, four CPP sites in Ohio, North Carolina, and New Hampshire, and two Brownfield sites in Florida and California rank among the most promising locations. These results underscore the potential of integrating machine learning and optimization techniques to transform nuclear siting, paving the way for a cost-effective and sustainable energy future.
@@
Nuclear Power Plants, Brownfield Sites, Site Selection, Multi-Objective Optimization, Coal to Nuclear Transition Word Count: 11419
journal: Energy Conversion and Management:X (published in 10.1016/j.ecmx.2025.100923)
@@
Feedforward Neural Network
@@
Highlights
  • 1.
    Analysis of the suitability of 30,000+ United States Brownfield and coal sites for hosting nuclear power plants.
  • 2.
    Multi-objective combinatorial methodology for nuclear power plant site ranking is developed.
  • 3.
    The method does not require analyst-defined weights in site objectives that could introduce bias.
  • 4.
    Data-driven neural network model is developed to predict site metrics for any location in the United States.
  • 5.
    Coal sites prove to be competitive for new nuclear plants, with some Brownfield sites also being feasible.
@@ 1 Introduction
Nuclear energy has several important applications, ranging from electricity generation to specialized industrial and medical uses [1, 2, 3, 4]. Nuclear energy ranks among the highest-yielding energy sources and has very low carbon dioxide emissions, even when accounting for the entire lifespan of the nuclear power plant (NPP). With its significant potential, nuclear energy could play a key role in achieving net-zero carbon emissions. However, a major challenge is its high levelized cost of electricity. While production costs for renewable energy sources have consistently declined over time, nuclear energy had the highest levelized cost of electricity in 2023, driven by rising capital expenditures and regulatory requirements [5]. In contrast, wind and solar energy tend to reduce electricity prices in their regions since their levelized costs are generally lower than the average for all energy sources [6]. Although the raw material costs for nuclear electricity production have also decreased [7], the substantial initial costs of nuclear power facilities remain a primary factor in nuclear energy’s higher prices.
In addition to economic challenges, socioeconomic and regulatory factors significantly influence nuclear energy adoption. Public sentiment, political support, and regulatory hurdles shape perceptions of nuclear energy and its feasibility as a sustainable energy solution [8, 9]. These factors also impact the repurposing of coal-fired power plants (CPPs) to Brownfield sites for nuclear energy projects, where public trust, local job creation, and community engagement play critical roles in project acceptance and success [10, 11]. Without addressing these challenges through comprehensive policies and transparent communication, the adoption of nuclear energy and the redevelopment of CPP sites for nuclear projects may face considerable resistance [8]. Furthermore, concerns around radioactive waste, safety, and infrastructure costs remain central to public discourse and influence both nuclear energy deployment and the feasibility of transforming CPPs into Brownfield nuclear sites [11, 12]. In order to increase public acceptance of NPPs and ensure public cooperation, as well as to avoid public backlash, the site of the CPP transformed into an NPP should be selected carefully, with public sentiment toward nuclear energy taken into consideration.
@@
By repurposing established facilities, the capital costs associated with NPP construction could be reduced. This analysis focuses on the United States, where we consider two existing facility options for new nuclear siting: coal power plant (CPP)s and Brownfields. First, we consider using the CPP station sites for NPP development. CPPs include control buildings, storage facilities, auxiliary structures, roads, electrical and water infrastructure, and substations. The presence of these existing facilities reduces the need for new construction, thereby lowering the capital costs associated with NPP installation. In comparison to a greenfield NPP building operation, the overall capital costs could be decreased by 15-35%, depending on the magnitude and choice of technology used in repurposing the framework of a coal power facility with a 1,200 MWe power output [13]. However, these CPP locations still have drawbacks. For example, if site remediation of a CPP is required before beginning the NPP construction project, the coal to nuclear repurposing costs could be high. Still, the existing educated workforce of the CPP would decrease the training and education costs of the personnel [14]. A complete cost analysis must be carried out before such a repurposing operation. According to the U.S. Department of Energy, 237 operational and 157 idle potential CPP sites are appropriate for hosting nuclear energy facilities [13]. Before reusing the CPP buildings, each of these sites must be thoroughly studied, as NPPs are complex, long-term projects with several risk factors and consequences. The NPP sitting in U.S. CPP locations has been researched in the past [15, 16, 17, 18]. CPP locations promise to be valuable sites for NPP siting. However, the number of U.S. CPPs is not enough to achieve the ultimate net-zero goals for the country. Therefore, an analysis that includes a more comprehensive look at all potential NPP sites in the U.S. is needed.
@@
The second option to host a NPP is to consider Brownfield sites. The term ”Brownfield” refers to formerly built land that may be contaminated, underutilized, or abandoned, and may occasionally have low pollution levels [19]. The economic costs and risks of Brownfield redevelopment projects are lower in areas that have an increased demand for residential and commercial development [20]. According to the U.S. Environmental Protection Agency (EPA), Assessment Cleanup and Redevelopment Exchange System (ACRES), Brownfield sites are properties where the existence or potential presence of a hazardous material, pollutant, or contaminant may make it more difficult to expand, redevelop, or reuse [21]. Common types of Brownfield sites are former industrial or manufacturing plants, abandoned gas stations, landfills or waste disposal areas, rail yards, mines, and junkyards. Repurposing these sites could benefith both the environment and the economy. To assess, clean up, and sustainably reuse the Brownfield sites, the EPA Brownfields Program provides grants and technical assistance to communities.
@@
The siting problem of NPP location has been an ongoing topic for the last 80 years. The problem consists of many different degrees of freedom and each requirement must be analyzed carefully since NPPs are long-term and high-cost projects. Most of the siting analyses in this field use a technique called the “weighted sum method”. In this method, explicit weights are applied to the geographical data to perform location selection studies for NPPs. To form an objective function, these techniques use pre-determined weights that the analyst justifies based on several factors. Geographical characteristics like seismicity, the presence of cooling water, fault lines, the distance to centers of population, international borders, and logistic network are the most crucial factors taken into account in this weighted sum method [16]. In some studies, only the most crucial characteristics are taken into account, such as seismic activity, population density, cooling water cost, reactor unit cost, and consumer proximity [17]. More sophisticated approaches use detailed NPP socioeconomic data. These methods handle financial, site-related, welfare, and project life-cycle factors independently in separate subroutines, by using their built-in socioeconomic data and weights [18]. Both simple and complex approaches still depend on users to assess the weights assigned to each site objective or attribute. The analyst responsible for setting thresholds, assigning weights, and processing raw data ultimately makes these critical decisions. In this process, the choice of weights can be subjective or objective, but never entirely removes the analyst’s bias in prioritizing certain sites for location selection. This is demonstrated in a recent study on CPP assessment in the U.S. for NPP transition conducted by [22], which involved assigning weights to various site objectives. While we recognize that the authors determined these weights through thorough research and extensive consultations, this level of rigor cannot be guaranteed in every study. Additionally, the approach still introduces bias, as site rankings inherently depend on the assigned weights. Therefore, an alternative method that avoids predetermined weights is needed to better identify ideal locations for NPP siting with limited bias. Multi-objective optimization techniques may offer some solution pathways to these challenges.
@@
A range of optimization methods for minimization and maximization problems are available in the literature, typically grouped into four main categories: stochastic, genetic, first-order, and second-order algorithms [23]. Most optimization algorithms can solve single-objective optimization problems when the objective function (also known as fitness or cost function) is a scalar. However, when addressing problems that involve multiple objectives, multi-objective optimization (MOO) approaches are required. These algorithms are designed to handle the complexity of balancing multiple objectives simultaneously during the search.
@@
In the problem of site selection, there is a high likelihood that multiple sites may have favorable characteristics which makes optimization challenging. We propose the concept of site characteristic “domination” as a solution to this problem. The method of non-dominated sorting is used in different optimization algorithms to identify optimal solutions in [24, 25, 26]. The most optimal samples in a non-dominated sorting operation are also known as the Pareto front, where each solution represents a trade-off across multiple objectives. In this Pareto-efficient set, no single solution is superior across all objectives, but the Pareto solutions that balance all objectives can be further analyzed. Consequently, in order to find the best NPP siting locations independent of any bias, we conduct a combinatorial analysis of Pareto front solutions by running a massive number of site attribute combinations, thereby mitigating external bias towards any site. This approach enables the identification of a ranked set of locations and the selection of the best site, rather than yielding inconclusive results in very high dimensional pareto fronts.
This research presents a new approach to nuclear reactor site assessment and Brownfield redevelopment by specifically targeting the siting of nuclear power plants NPPs; an area that has not been thoroughly explored in the existing literature. Utilizing raw site data available through the Siting Tool for Advanced Nuclear Development (STAND) tool [27], we created a dataset of Brownfield NPP siting locations, focusing on 34,211 sites that do not require extensive industrial cleanup, thereby establishing a robust resource for future nuclear siting projects. Furthermore, we conducted a comparative analysis between CPPs and Brownfields, challenging the prevailing assumption that CPP sites are the only sites that offer optimal conditions for NPP siting. By comparing 265 CPPs with the Brownfield sites, our findings illuminate the feasibility of alternative sites, thereby offering additional compelling options to the current trends in coal-to-nuclear transition research. To enhance the efficiency of future site assessment, we developed a machine learning model that predicts siting metrics and objective importance; significantly reducing the computational burden associated with traditional methodologies. This model is poised to facilitate rapid assessments for any contiguous U.S. location, ultimately streamlining the decision-making process in NPP siting. Accordingly, the key contributions of this study can be summarized as:
@@ 2 Data Collection and Processing
In the Brownfield nuclear reactor siting analysis, we use the ACRES Brownfield database (described above) as a starting point. We gathered the geographical database files from the EPA. We extracted the coordinates recorded in ACRES from the database layers. We then use the extracted coordinates in the Siting Tool for Advanced Nuclear Development (STAND) tool [27] developed by the National Reactor Innovation Center at Idaho National Laboratory and maintained by the Fastest Path to Zero Initiative at the University of Michigan [28]. We processed each extracted Brownfield coordinate in STAND to creating a Brownfield nuclear reactor siting dataset that includes all Brownfield sites with their corresponding coordinates and site characteristics. We found unreliable data for Alaska and Hawaii in the STAND tool, because of data availability issues; the Brownfield sites in these two states were excluded from the final dataset. The ACRES U.S. Brownfield locations gathered from EPA are given in Figure 1. The Brownfield siting dataset is current as of May 2024.
About 34,211 different contiguous U.S. Brownfield coordinates in the EPA ACRES Brownfield database have been processed through STAND to generate values for 22 unique objectives. The attributes in the dataset describe the characteristics of each site. We chose these objectives to match the 1000 MWt or higher power reactor siting parameters used in the coal-to-nuclear transition analysis performed by the previous NPP siting research [22]. We classify the objectives selected for the Brownfield sites under three categories:
@@
2.1 Rationale for Objective Selection
The STAND tool identifies 22 critical socioeconomic, safety, and proximity characteristics as key attributes for nuclear power plant (NPP) siting [27]. These characteristics, selected for their relevance, are justified in the Site Comparison step within the Relevance Form tab of the tool. The selection of these attributes reflects a comprehensive understanding of technical feasibility, safety imperatives, and socioeconomic considerations, ensuring that the siting process aligns with both regulatory requirements and public expectations. By integrating these diverse criteria, the STAND tool facilitates informed decision-making for sustainable and publicly acceptable nuclear energy development [29, 30].
Several socioeconomic attributes significantly influence NPP siting. State-level restrictions present a key challenge, ranging from procedural barriers to outright moratoria, requiring substantial time and resources to address [30]. Retail electricity prices act as a critical indicator of nuclear energy’s cost competitiveness in specific states [31]. States with energy consumption exceeding their production are more inclined to prioritize new energy facilities, including NPPs, to ensure energy security [31]. Public attitude toward nuclear energy is essential for community engagement and the success of siting projects [32]. Additionally, market regulation affects factors such as capital costs, financing, and industry partnerships, which shape the feasibility of NPP siting [29]. Local labor rates further influence the economics of construction and maintenance [31]. Finally, the inclusion of the Social Vulnerability Index (SVI) ensures equitable site selection by addressing risks to vulnerable populations and promoting resilience [33].
Safety is paramount in NPP siting, requiring thorough evaluations of hazardous facilities, seismic risks, and other geotechnical factors. Proximity to industrial sites, such as chemical plants or refineries, introduces risks including toxic releases or flammable materials, necessitating careful assessment [30]. Seismic considerations, such as fault lines within a 200-mile radius and ground motion potential, significantly impact site suitability and design requirements [34]. Site-specific factors like liquefaction, landslides, and subsidence are critical for ensuring the integrity of safety-related functions [29]. Wetlands and floodplains are evaluated under regulations like Executive Order 11988 and the Clean Water Act to avoid ecological damage and mitigate flooding risks [30]. Additionally, steep slopes are avoided due to the high economic costs and technical challenges associated with construction [35].
Proximity considerations ensure NPPs are located in areas conducive to operational efficiency and minimal risk. Regulatory guidelines, such as U.S. Nuclear Regulatory Commission (NRC) Regulatory Guide 4.7, recommend population density thresholds to ensure emergency preparedness and safety [34]. Reusing infrastructure from closed facilities can streamline development and reduce costs [30]. Proximity to nuclear research facilities enhances technical support, while access to electrical substations ensures seamless integration with the power grid [29]. Transportation infrastructure, including heavy-haul routes, is critical for delivering modular reactor components [30]. Finally, freshwater availability is essential for reactor cooling, making proximity to adequate water resources a key criterion for site selection [30].
The STAND tool provides a holistic framework for NPP siting, by addressing these socioeconomic, safety, and proximity factors. This approach balances technical feasibility with safety imperatives and socioeconomic considerations, fostering public acceptance and compliance with regulatory requirements. The integration of these criteria into our analysis enhances the alignment of nuclear energy development with sustainability goals and community resilience, thereby strengthening the role of nuclear power in the transition to clean energy [29, 30, 31, 34].
Figure 1: ACRES Brownfield locations in the United States provided by the Environmental Protection Agency.